OpenStudy (anonymous):

What is the implicit derivitive of: 3x²y+9y²x=-6

OpenStudy (abb0t):

For every y use y' or dy/dx. \[3x^2\frac{ dy }{ dx } + 6xy + 9y^2 + 18y \frac{ dy }{ dx }x = 0\] Use algebra to rearrange and solve for dy/dx.

OpenStudy (anonymous):

Oh ok, I had the right idea but it looks like I was backwards.

OpenStudy (abb0t):

Your answer should look something like this: \[\frac{ dy }{ dx } = function\] or \[y ' = function\]

OpenStudy (anonymous):

Ok yeah I was putting y' in the wrong places. Thank you for the help.